Chapter 19, Problem 13 Two point charges, +3.10 HC and -6.60 uC, are separated by 2.80 m. What is the electric potential midway between them? Number the tolerance is +/-5% Units

Chapter 19, Problem 24 Two identical point charges (q-+1.30 x 10。C) are fixed at opposite corners of a square whose sides have a length of 0.540 m. A test charge (a)- 9.60 x 10 C 1.40 x 10 kg, is released from rest at one of the corners of the square. Determine the speed of the test charge when it reaches the center of the square. with a mass of 70 UnitsY Number the tolerance is +/-590
Answer #1

electric potential due to charge = kq/r

where q = charge , r =distance from the charge

electric potential midway = kq1/r1 + kq2/r2

= 9*10^9*10^-6 ( 3.1 / 1.4 - 6.6 / 1.4)

= -22500 V